Brazil: Balance of Payment

                                                                                                Unit: millions of USDs

Current Account

Credits

Debits

Exports

41,971

            Goods

28,113

            Services

5,028

            Income

8,830

Imports

? (1)

            Goods

-7,892

            Services

-35,835

            Income

-48,397

Unilateral transfer

? (2)

-186

Trade Balance

? (3)

Balance on current account

-47,365

Financial Account

Direct Investment

48,438

-11,519

Portfolio Investment

74,006

-10,995

Other investment

1,119

-1,387

Balance on financial account

? (4)

Statistical discrepancies

? (5)

Official Reserve Account

-49,101

A) What are the Brazil imports in Current Account?

b. -92,124

d. -50,153

a. -41,971

c. -92,310

B) What is the Brazil transfer payments to foreign entities (credit side of Unilateral Transfer)?

a. 186

b. 2,974

d. 50,339

c. 3,790

C) What is the Brazil Trade Balance?

d. 5,394

b. 10,586

c. 20,221

a. 50,153

D) What is the Brazil balance on financial account?

a. 50,561

d. 123,563

c. 99,662

b. 74,462

E) What is the Brazil Statistical discrepancies?

c. -3,196

a. -49,101

b. -23,901

d. -25,200
Answer #1

Answer

(A) b. -92124

Brazil imports in current account = Goods + services + Income

= -7892-35835-48397

= -92124

(B) b.2974

Given balance on current account = -47365

Total exports -Total imports = Balance on current account

41971 + unilateral transfer - (92124+186) = -47365

therefore unilateral transfer balance credit side = 2974

(C) c.20221

Brazil trade balance = Export of goods - Import of goods

= 28113-7892 = 20221

(D) c.99662

Brazil balance on financial account = 48438+74006+1119-11519-10995-1387

= 99662

(E) c.-3196

Brazil statistical discrepancies = -47365-49101+99662

= -3196